Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2018-08-31

Observers: M. Gromadzki, T. Wevers
Support Team: G. Pignata, O. Rodriguez
TAT: L. Galbany, C. Barbarino

Calibrations: 
EFOSC bias+calib, SOFI calibration afternoon

Weather: 
thick clouds, dome closed until UT6:30, then it was quite humid (45%) and seeing 1.1-1.9 arcsec

UT       Target       Type      Setup          Exp         airmass  DIMM  note
-----------------------------------------------------------------------------------------
                                             
EFOSC                                                                       
06:35    AT2018fit    Class     gr13/1.5"      900          1.44    1.76   SNII  z=0.01
06:56    AT2018fms    Class     gr13/1.5"      600          1.3     1.9    SNIIb z=0.6
07:16    AT2018fka    Class                                 1.3     1.9    image analysis - object to faint
07:30    AT2018fmt    Class     gr13/1"        900          1.04    1.4    blue continuum faint narrow Balmers z=0.054
07:48    AT2018ffo    Class     gr13/1.5"      1200         1.27    1.2    SNIa +40d z(SNID)=0.040      
08:12    AT2018fhx    Class     gr13/1"        300          1.7     1.4    SNIa +5d  z(SNID)=0.027
08:33    SN2018fds    Follow    gr11/1"        1400+1800    1.02    1.1    First exposure interrupted  
09:40    VMA2         STD    gr11/13/16                     1.7     1.5




Comments